# StylizedGS: Controllable Stylization for 3D Gaussian Splatting

### v2

@@ Abstract
As XR technology continues to advance rapidly, 3D generation and editing are increasingly crucial. Among these, stylization plays a key role in enhancing the appearance of 3D models. By utilizing stylization, users can achieve consistent artistic effects in 3D editing using a single reference style image, making it a user-friendly editing method. However, recent NeRF-based 3D stylization methods encounter efficiency issues that impact the user experience, and their implicit nature limits their ability to accurately transfer geometric pattern styles. Additionally, the ability for artists to apply flexible control over stylized scenes is considered highly desirable to foster an environment conducive to creative exploration. To address the above issues, we introduce StylizedGS, an efficient 3D neural style transfer framework with adaptable control over perceptual factors based on 3D Gaussian Splatting (3DGS) representation. We propose a filter-based refinement to eliminate floaters that affect the stylization effects in the scene reconstruction process. The nearest neighbor-based style loss is introduced to achieve stylization by fine-tuning the geometry and color parameters of 3DGS, while a depth preservation loss with other regularizations is proposed to prevent the tampering of geometry content. Moreover, facilitated by specially designed losses, StylizedGS enables users to control color, stylized scale, and regions during the stylization to possess customization capabilities. Our method achieves high-quality stylization results characterized by faithful brushstrokes and geometric consistency with flexible controls. Extensive experiments across various scenes and styles demonstrate the effectiveness and efficiency of our method concerning both stylization quality and inference speed.
@@ 1 Introduction
In this paper, we introduce the first controllable scene stylization method based on 3DGS, StylizedGS. Given a reference style image, our method effectively transfer the style features to the entire 3D scene represented by a set of 3D Gaussians. Our method facilitates the artistic creation of visually coherent novel views that exhibit transformed and detailed style features in a visually reasonable manner. More importantly, StylizedGS operates at a fast inference speed, ensuring efficiency when rendering stylized scenes. To achieve effective stylization and avoid geometry distortion, a straightforward approach is to fine-tune the color of 3D Gaussians, minimizing both style loss and content loss. However, focusing solely on color fails to effectively capture the overall stylistic features when the style patterns contain intricate details. Instead, we propose formulating the 3DGS stylization as a joint optimization of both the geometry and color of 3D Gaussians to capture the detailed style features while preserving the semantic content. Specifically, our method employs a two-stage stylization framework. In the first stage, we optimize the 3D Gaussians to align the color distribution with the style image while reducing artifacts and geometry noises using our filter-based 3D Gaussian refinement. In the second stage, we achieve final stylization by applying a nearest neighbor feature match (NNFM) loss, optimizing the 3D Gaussians’ parameters to capture coherent fine-grained style details.
To enable users to customize the stylization process, we propose a series of strategies and loss functions to control perceptual factors such as color, scale, and spatial regions  [24, 25, 26, 27] in the final stylized results. By incorporating this enhanced level of perceptual controllability, our approach synthesizes diverse 3D scene stylization results with varied visual characteristics. Furthermore, we introduce a depth preservation loss that eliminates the necessity for additional networks or regularization operations to preserve the learned 3D scene geometry. It establishes a delicate balance between geometric optimization and depth preservation, facilitating effective style pattern transfer while mitigating significant deterioration of geometric content.
@@ 2 Related Work
3 Method
Given a 3D scene represented by a collection of captured images with corresponding camera parameters, our objective is to efficiently accomplish consistent style transfer from an arbitrary 2D style image to the 3D scene. Building on recent advances in 3D Gaussian Splatting (3DGS) [17], we optimize the 3DGS representation to generate stylized scenes with intricate artistic features and high visual quality. Our stylization process consists of two key steps: 1) We first recolor the 3D scene to align its color statistics with those of the style image. Simultaneously, we adopt a filter-based 3D Gaussian refinement scheme to minimize the impact of floaters in the reconstruction, which is crucial for ensuring the final stylization quality. 2) We then optimize 3D Gaussians further by exploiting nearest-neighbor feature matching style loss to capture detailed local style patterns. During the optimization, we incorporate a depth preservation loss and regularization terms to preserve the geometric features of the scene. Finally, we introduce how to achieve flexible control over the color, scale, and spatial attributes in our 3DGS stylization (Sec. 3.3). As a result, users can create stylized renderings with customized artistic expression, and explore the consistently stylized scene in a free-view manner. The pipeline of our method is shown in Fig. 2.
3.1 Preliminaries: 3D Gaussian Splatting
Gaussian Splatting [17] encapsulates 3D scene information using a set of 3D colored Gaussians. This technique exhibits rapid inference speeds and exceptional reconstruction quality compared to NeRF. To represent the scene, each Gaussian is described by a centroid $p=\{x,y,z\}\in\mathbb{R}^{3}$, a 3D vector $s\in\mathbb{R}^{3}$ for scaling, and a quaternion $q\in\mathbb{R}^{4}$ for rotation. Additionally, an opacity value $\alpha\in\mathbb{R}$ and a color vector $c$ represented in the coefficients of a spherical harmonic (SH) function of degree 3 are used for fast alpha-blending during rendering. These trainable parameters are collectively symbolized by $G_{\theta_{i}}$, where $G_{\theta_{i}}=\{p_{i},s_{i},q_{i},\alpha_{i},c_{i}\}$, representing the parameters for the $i$-th Gaussian. To visualize the 3D Gaussians and supervise their optimization, 3DGS projects them onto the 2D image plane. The implementation leverages differentiable rendering and gradient-based optimization on each pixel for the involved Gaussians. The pixel color $c^{\alpha}$ is determined by blending the colors $c_{i}$ of those ordered Gaussians that overlap the pixel. This process can be formulated as:
@@ 3.1 Preliminaries: 3D Gaussian Splatting
3.2 Style Transfer to 3D Gaussian Splatting
@@
Filter-based Refinement. After the color transfer step, the color attributes $c^{re}$ of 3DGS and the recolored rendering images $\mathcal{I}_{content}^{re}$ are consistent with the palette of the style image. However, some floaters in the original 3DGS will also be colored during this process, significantly affecting the quality of the stylization, as shown in Fig. 13. These colored floaters often result from Gaussians with excessively large scales or improper opacity values. To address this issue, we design a filter-based refinement process to exclude these floaters while ensuring reconstruction quality before stylization. We use the recolored content images $\mathcal{I}_{content}^{re}$ as supervision to fine-tune the 3DGS with the recolored color parameter $c^{re}$ and filter the 3D Gaussian floater at every certain number of fine-tuning iterations. The filtering is performed based on the scale and opacity of the 3D Gaussians. Specifically, Gaussians whose sizes are in the top kor whose opacities are in the lowest kwill be filtered out. The threshold kremains constant during the fine-tuning process. By maintaining a fixed threshold, we ensure that the reconstructed scene does not change significantly, as the fine-tuning process is relatively short. Such a refinement step minimizes the impact on the overall scene reconstruction quality and enhances color correspondence, while the GS filter removes the floaters. During fine-tuning, we incorporate the reconstruction loss from [17]:
@@
where, $(i*,j*)=\arg\min_{i^{\prime},j^{\prime}}D(F_{render}(i,j),F_{style}(i^{\prime},j^{\prime}))$, and
$D(a,b)$ is the cosine distance between two feature vectors $a$ and $b$. To preserve the original content structure during the stylization, we additionally minimize the mean squared distance between the content feature $F_{content}$ and the stylized feature $F_{render}$:
$\mathcal{L}_{content}=\frac{1}{H\times W}||F_{content}-F_{render}||^{2}$ (5)
where $H$ and $W$ represent the size of rendered images. While the content loss $\mathcal{L}_{content}$ mitigates the risk of excessive stylization, optimizations applied to the geometric parameters of 3D Gaussians can still induce changes in scene geometry. To address this, we introduce a depth preservation loss to maintain general geometric consistency without relying on any additional depth estimator. It should be noticed that there is a trade-off between geometric optimization and depth preservation loss, allowing our stylization to effectively transfer style patterns while avoiding potential damage to geometric content. Specifically, in the color transfer stage, we initially generate the original depth map $D_{origin}$ by employing the alpha-blending method of 3DGS. The alpha-blended depth $d^{\alpha}$ of each pixel is calculated as $d^{\alpha}=\sum_{i}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value for $i$-th Gaussian. During the stylization, we minimize the $L_{2}$ loss between the rendered depth map $D_{render}$ and the original one $D_{origin}$:
$\mathcal{L}_{depth}=\frac{1}{H\times W}||D_{origin}-D_{render}||^{2}$ (6)
@@
$\mathcal{L}_{reg}^{ds}=\frac{1}{M}||\Delta s||,\quad\mathcal{L}_{reg}^{d\alpha}=\frac{1}{M}||\Delta\alpha||$ (7)
@@
3.3 Perceptual Control
Building on the previous components of our approach, we can achieve both rapid and intricate style transfer while preserving the original content. Here, we address the challenge users face in customizing the stylization process. Inspired by [24], we propose a series of strategies and loss functions for 3DGS stylization to control various perceptual factors, including color, scale, and spatial areas, enhancing the customization capabilities of our approach. The stylization loss will be replaced with a different form, namely $L_{style}^{color}$, $L_{style}^{scale}$ or $L_{style}^{spatial}$ defined below.
Color Control. Color information within an image is a crucial perceptual aspect of its style. Unlike other stylistic elements such as brush strokes or dominant geometric shapes, color is largely independent and stands out as a distinct characteristic. As illustrated in Fig. 3.2, users may seek to preserve the original color scheme by transferring only the pattern style or by combining the pattern style of one image with the color style of another. Consequently, our method provides users with independent control over color style during the stylization process. To achieve this, in our color transfer step, we can pre-color 3D scenes to align with any user-desired hues. Similar to the luminance-only algorithm proposed in [56], we utilize the YIQ color space to separate the learning of luminance and color information. As defined in the color style loss $\mathcal{L}_{style}^{color}$ in Eq. 9, we extract the luminance channel from the rendered view $\mathcal{I}_{render}^{Y}$ and the style image $\mathcal{I}_{style}^{Y}$ for the luminance-only style loss calculation. Meanwhile, the RGB channels are retained for the content loss $\mathcal{L}_{content}$ calculation.
@@ 3.4 User Interface
4 Experiments
@@
Datasets. We conduct extensive experiments on multiple real-world scenes inLLFF [59] which contains forward-facing scenes and Tanks & Temples (T&T) dataset [60] which includes unbounded 360 large outdoor scenes. Furthermore, our experiments involve a diverse set of style images from [13] and WikiArt [61], allowing us to assess our method’s capability to handle a wide range of stylistic exemplars.
@@
4.1 Qualitative Comparisons
We show visual comparisons on the LLFF dataset in Fig. 7. By jointly optimizing the opacity and color components, our method performs better in learning intricate geometric patterns and meso-structures. Specifically, for the scene of “trex” with a Lego style, other methods capture only the color style or a basic square pattern, whereas our approach successfully illustrates the detailed structure of the Lego blocks, clearly depicting even the small cylindrical patterns. This is also evident in the scene of ”horns” with a snow style, where our method produces visually more pleasing results that faithfully match the style of the hexagonal snowflake structure. Additionally, our method preserves more semantic information in the stylized scene by utilizing the depth map to maintain spatial distribution in the image. We show more stylization results on the LLFF dataset in Fig. 9.
Fig. 8 presents comparison results on the T&T dataset. It can be seen that our method achieves a better style match with the style image compared to the others. For instance, LSNV tends to produce overly smoothed results and lacks intricate structures, while ARF struggles to consistently transfer the style across the entire scene and exhibits deficiencies in color correspondence, such as the results of the second group. StyleGaussian does not faithfully transfer the style, resulting in blurry results. In contrast, our method accurately captures both the color tones and brushstrokes throughout the entire scene, demonstrating superior style matching compared to the baselines. We show more stylization results on the T&T dataset in Fig. 10.
TABLE I: Quantitative comparisons on stylization under novel views. We report ArtFID, SSIM, DISTS, average training time (Avg. train), and average rendering FPS (Avg. FPS) for our method and other baselines. ‘-’ denotes that the method does not require individual stylization training.
Metrics ArtFID($\downarrow$) SSIM($\uparrow$) DISTS($\downarrow$) Avg. train($\downarrow$) Avg. FPS($\uparrow$)
@@ 4.1 Qualitative Comparisons
4.2 Quantitative Comparisons
A significant advantage of the 3DGS representation is the capability of rendering high-quality views in real-time and with much less training time. Therefore, we conducted quantitative comparisons with other methods to demonstrate both the high fidelity and efficiency of our approach. As the optimization time during the stylization process is notably influenced by the resolution, we randomly selected 50 stylized scenes with consistent resolution from the LLFF dataset and evaluated the average times. For the stylization metrics, we choose ArtFID [63] to measure the quality of stylization following [64]. We also use SSIM [65] and DISTS [66] between the content and stylized images to measure the performance of detail preservation and structural similarity following [35, 67, 68, 69, 70]. As shown in Tab. I, our method excels in both stylized visual quality and content preservation. Regarding efficiency, our method has significant advantages in average rendering FPS (Avg. FPS) and can achieve real-time free-view synthesis. At the same time, it outperforms other optimization-based stylization methods in terms of average training time (Avg. train) for stylization and can achieve per-scene stylization within 1 minute on a single NVIDIA RTX 4090 GPU. Note that while StyleRF and StyleGaussian are feed-forward methods and do not require individual stylization training for each case, it’s worth mentioning that StyleRF’s rendering is slow and cannot be viewed in real time.

### v3

@@ Abstract
As XR technology continues to advance rapidly, 3D generation and editing are increasingly crucial. Among these, stylization plays a key role in enhancing the appearance of 3D models. By utilizing stylization, users can achieve consistent artistic effects in 3D editing using a single reference style image, making it a user-friendly editing method. However, recent NeRF-based 3D stylization methods encounter efficiency issues that impact the user experience, and their implicit nature limits their ability to accurately transfer geometric pattern styles. Additionally, the ability for artists to apply flexible control over stylized scenes is considered highly desirable to foster an environment conducive to creative exploration. To address the above issues, we introduce StylizedGS, an efficient 3D neural style transfer framework with adaptable control over perceptual factors based on 3D Gaussian Splatting representation. We propose a filter-based refinement to eliminate floaters that affect the stylization effects in the scene reconstruction process. The nearest neighbor-based style loss is introduced to achieve stylization by fine-tuning the geometry and color parameters of 3DGS, while a depth preservation loss with other regularizations is proposed to prevent the tampering of geometry content. Moreover, facilitated by specially designed losses, StylizedGS enables users to control color, stylized scale, and regions during the stylization to possess customization capabilities. Our method achieves high-quality stylization results characterized by faithful brushstrokes and geometric consistency with flexible controls. Extensive experiments across various scenes and styles demonstrate the effectiveness and efficiency of our method concerning both stylization quality and inference speed.
@@ I Introduction
In this paper, we introduce the first controllable scene stylization method based on 3DGS, StylizedGS. Given a reference style image, our method effectively transfer the style features to the entire 3D scene represented by a set of 3D Gaussians. Our method facilitates the artistic creation of visually coherent novel views that exhibit transformed and detailed style features in a visually reasonable manner. More importantly, StylizedGS operates at a fast inference speed, ensuring efficiency when rendering stylized scenes. To achieve effective stylization and avoid geometry distortion, a straightforward approach is to fine-tune the color of 3D Gaussians, minimizing both style loss and content loss. However, focusing solely on color fails to effectively capture the overall stylistic features when the style patterns contain intricate details. Instead, we propose formulating the 3DGS stylization as a joint optimization of both the geometry and color of 3D Gaussians to capture the detailed style features while preserving the semantic content. Specifically, our method employs a two-stage stylization framework. In the first stage, we optimize the 3D Gaussians to align the color distribution with the style image while reducing artifacts and geometry noises using our filter-based 3D Gaussian refinement. In the second stage, we achieve final stylization by applying a nearest neighbor feature match (NNFM) loss, optimizing the 3D Gaussians’ parameters to capture coherent fine-grained style details. To enable users to customize the stylization process, we propose a series of strategies and loss functions to control perceptual factors such as color, scale, and spatial regions  [24, 25, 26, 27] in the final stylized results. By incorporating this enhanced level of perceptual controllability, our approach synthesizes diverse 3D scene stylization results with varied visual characteristics. Furthermore, we introduce a depth preservation loss that eliminates the necessity for additional networks or regularization operations to preserve the learned 3D scene geometry. It establishes a delicate balance between geometric optimization and depth preservation, facilitating effective style pattern transfer while mitigating significant deterioration of geometric content.
@@ II Related Work
III Method
Our method enables efficient 3D style transfer by optimizing 3DGS representations [17] to transform captured scenes into consistently stylized outputs using arbitrary 2D style images. The pipeline comprises two sequential optimization stages: First, global scene recoloring aligns color statistics with the style image while filtering reconstruction floaters through Gaussian refinement. Second, local texture detailing employs nearest-neighbor feature matching to transfer intricate artistic patterns, regularized by depth preservation to maintain geometric integrity. This framework provides flexible control over color, scale, and spatial attributes (Sec. III-C), enabling customizable artistic expression and consistent free-view rendering as demonstrated in Fig. 2.
III-A Preliminaries: 3D Gaussian Splatting
Gaussian Splatting [17] encapsulates 3D scene information using a set of 3D colored Gaussians. This technique exhibits rapid inference speeds and exceptional reconstruction quality compared to NeRF. To represent the scene, each Gaussian is described by a centroid $x\in\mathbb{R}^{3}$, a 3D vector $s\in\mathbb{R}^{3}$ for scaling, and a quaternion $q\in\mathbb{R}^{4}$ for rotation. Additionally, an opacity value $\alpha\in\mathbb{R}$ and a color vector $c$ represented in the coefficients of a spherical harmonic (SH) function of degree 3 are used for fast alpha-blending during rendering. These trainable parameters are collectively symbolized by $G_{\theta_{i}}$, where $G_{\theta_{i}}=\{x_{i},s_{i},q_{i},\alpha_{i},c_{i}\}$, representing the parameters for the $i$-th Gaussian. To visualize the 3D Gaussians and supervise their optimization, 3DGS projects them onto the 2D image plane. The implementation leverages differentiable rendering and gradient-based optimization on each pixel for the involved Gaussians. The pixel color $c^{\alpha}$ is determined by blending the colors $c_{i}$ of those ordered Gaussians that overlap the pixel. This process can be formulated as:
@@ III-A Preliminaries: 3D Gaussian Splatting
III-B Style Transfer to 3D Gaussian Splatting
@@
Filter-based Refinement. After the color transfer step, the color attributes $c^{re}$ of 3DGS and the recolored rendering images $\mathcal{I}_{content}^{re}$ are consistent with the palette of the style image. However, some floaters in the original 3DGS will also be colored during this process, significantly affecting the quality of the stylization, as shown in Fig. 17. To address this issue, we design a filter-based refinement process to exclude these floaters while ensuring reconstruction quality before stylization. We use the recolored content images $\mathcal{I}_{content}^{re}$ as supervision to fine-tune the 3DGS with the recolored color parameter $c^{re}$ and filter the 3D Gaussian floater at every certain number of fine-tuning iterations. The filtering is performed based on the scale and opacity of the 3D Gaussians. Specifically, Gaussians whose sizes are in the top k% or whose opacities are in the lowest k% will be filtered out. The threshold k% remains constant during the fine-tuning process. By maintaining a fixed threshold, we ensure that the reconstructed scene does not change significantly, as the fine-tuning process is relatively short. During fine-tuning, we incorporate the reconstruction loss from [17]:
@@
where, $(i*,j*)=\arg\min_{i^{\prime},j^{\prime}}D(F_{render}(i,j),F_{style}(i^{\prime},j^{\prime}))$, and $D(a,b)$ is the cosine distance between two feature vectors $a$ and $b$. To preserve the original content structure during the stylization, we additionally minimize the mean squared distance between the content feature $F_{content}$ and the stylized feature $F_{render}$:
[…]
where $H$ and $W$ represent the size of rendered images. While the content loss $\mathcal{L}_{content}$ mitigates the risk of excessive stylization, optimizations applied to the geometric parameters of 3D Gaussians can still induce changes in scene geometry. To address this, we introduce a depth preservation loss to maintain general geometric consistency without relying on any additional depth estimator. Specifically, in the color transfer stage, we initially generate the original depth map $D_{origin}$ by employing the alpha-blending method of 3DGS. The alpha-blended depth $d^{\alpha}$ of each pixel is calculated as $d^{\alpha}=\sum_{i}T_{i}\alpha_{i}d_{i}$, where $d_{i}$ is the depth value for $i$-th Gaussian. During the stylization, we minimize the $L_{2}$ loss between the rendered depth map $D_{render}$ and the original one $D_{origin}$:
[…]
@@
[…]
@@
III-C Perceptual Control
Building on the previous components of our approach, we can achieve both rapid and intricate style transfer while preserving the original content. Here, we address the challenge users face in customizing the stylization process. Inspired by [24], we propose a series of strategies and loss functions for 3DGS stylization to control various perceptual factors, including color, scale, and spatial areas. The stylization loss will be replaced with a different form, namely $L_{style}^{color}$, $L_{style}^{scale}$ or $L_{style}^{spatial}$ defined below.
Color Control. Color information within an image is a crucial perceptual aspect of its style. Unlike other stylistic elements such as brush strokes or dominant geometric shapes, color is largely independent and stands out as a distinct characteristic. As illustrated in Fig. 3, users may seek to preserve the original color scheme by transferring only the pattern style or by combining the pattern style of one image with the color style of another. To support this flexibility, our method offers independent control over color during stylization.
We achieve this by pre-coloring the 3D scene to match user-specified hues, followed by style transfer in the YIQ color space, which decouples luminance and chrominance. Specifically, our color style loss $\mathcal{L}{\text{style}}^{\text{color}}$ (Eq. 9) is computed using only the luminance channel of the rendered view $\mathcal{I}{\text{render}}^{Y}$ and the style image $\mathcal{I}{\text{style}}^{Y}$. Meanwhile, RGB channels are retained for the content loss $\mathcal{L}{\text{content}}$, allowing for structural fidelity to be maintained during stylization.
@@ III-D User Interface
IV Experiments
Figure 9: Qualitative comparisons with the baseline methods on MipNeRF-360 datasets. (Please zoom in for better view.)
@@ IV Experiments
Datasets. We conduct extensive experiments on multiple real-world scenes in LLFF [58] which contains forward-facing scenes, Tanks & Temples (T&T)  [59] and MipNeRF-360 [60] dataset which includes unbounded 360 large outdoor scenes. Additionally, we demonstrate the generality of our method by presenting stylization results on the Deep Blending [61] and Instruct-NeRF2NeRF [62] datasets. Furthermore, our experiments involve a diverse set of style images from [13] and WikiArt [63], allowing us to assess our method’s capability to handle a wide range of stylistic exemplars.
@@
IV-A Qualitative Comparisons
Figure 12: Stylization results on MipNeRF-360 dataset.
We show visual comparisons on the LLFF dataset in Fig. 7. By jointly optimizing the opacity and color components, our method performs better in learning intricate geometric patterns and meso-structures. Specifically, for the scene of “trex” with a Lego style, other methods capture only the color style or a basic square pattern, whereas our approach successfully illustrates the detailed structure of the Lego blocks, clearly depicting even the small cylindrical patterns. This is also evident in the scene of ”horns” with a snow style, where our method produces visually more pleasing results that faithfully match the style of the hexagonal snowflake structure. Additionally, our method preserves more semantic information in the stylized scene by utilizing the depth map to maintain spatial distribution in the image.
Fig. 8 and Fig. 9 presents comparison results on the T&T and MipNeRF-360 dataset. It can be seen that our method achieves a better style match with the style image compared to the others. For instance, LSNV tends to produce overly smoothed results and lacks intricate structures, while ARF and Ref-NPR struggles to consistently transfer the style across the entire scene and exhibits deficiencies in color correspondence, such as the results of the second group. StyleGaussian does not faithfully transfer the style, resulting in blurry and noise results. In contrast, our method accurately captures both the color tones and brushstrokes throughout the entire scene, demonstrating superior style matching compared to the baselines.
We present additional stylization results across different datasets: in Fig. 10, Fig. 11, Fig. 13 and Fig. 14. Please refer to our supplementary for more stylization results.
Figure 13: Stylization results on Instruct-NeRF2NeRF dataset.
Figure 14: Stylization results on Deep Blending dataset.
TABLE I: Quantitative comparisons on stylization under novel views. We report ArtFID, SSIM, DISTS, average training time (Train), and average rendering FPS (FPS) for our method and other baselines. ‘-’ denotes that the method does not require individual stylization training.
Metrics ArtFID($\downarrow$) SSIM($\uparrow$) DISTS($\downarrow$) Train($\downarrow$) FPS($\uparrow$)
@@ IV-A Qualitative Comparisons
TABLE II: Quantitative comparisons on stylization under novel views on MipNeRF-360 dataset. We report ArtFID, SSIM, DISTS and average rendering FPS (FPS) for our method and other baselines.
Metrics ArtFID ($\downarrow$) SSIM ($\uparrow$) DISTS ($\downarrow$) FPS($\uparrow$)
ARF 34.79 0.25 0.35 7.6
Ref-NPR 35.16 0.27 0.33 5.3
StyleGaussian 45.38 0.24 0.32 116
Ours 29.12 0.40 0.28 137
TABLE III: Quantitative results. We evaluate the performance of our method against the state-of-the-art in terms of consistency, using LPIPS ($\downarrow$) and RMSE ($\downarrow$).
Methods
Short-range
Consistency
Long-range
Consistency
LPIPS RMSE LPIPS RMSE
ARF 0.036 0.051 0.185 0.128
Ref-NPR 0.035 0.052 0.173 0.124
StyleGaussian 0.031 0.048 0.159 0.117
Ours 0.028 0.044 0.112 0.088
IV-B Quantitative Comparisons
To assess both fidelity and efficiency, we perform quantitative comparisons with existing methods. For the stylization metrics, we choose ArtFID [65] to measure the quality of stylization following [66]. We also use SSIM [67] and DISTS [68] between the content and stylized images to measure the performance of detail preservation and structural similarity following [35, 69, 70, 71, 72]. As shown in Tab. I and Tab. II, our method excels in both stylized visual quality and content preservation. In terms of efficiency, our approach supports real-time free-viewpoint rendering and significantly reduces per-scene stylization time, achieving sub-minute training on a single RTX 4090 thanks to 3DGS. While StyleRF and StyleGaussian are feed-forward methods that require no per-scene training, they suffer from slower rendering and reduced stylization fidelity.
We perform quantitative comparisons on multi-view consistency in Tab. III, adopting established view-warping protocols [9, 11, 7, 64, 22, 73]. Using optical flow [74] with softmax splatting [75], we warp stylized views between camera poses and compute masked RMSE and LPIPS [76] scores to quantify consistency. Our evaluation spans 5 unbounded scenes from T&T and MipNeRF-360 datasets, employing 20 view pairs per scene at short-range (5-frame gap) and long-range (35-frame gap) intervals. Each pair is stylized with 10 distinct style images, yielding 200 comprehensive test cases. As evidenced in Tab. III, our method demonstrates superior consistency metrics, overcoming the view-dependent artifacts inherent in NeRF-based approaches and the feature decoding limitations of StyleGaussian’s 3D CNN architecture.